A square is inscribed in a circle. If the area of the square is 9 in^2, what is the ratio of the circumference of the circle to the area of the circle?
Answer by Alan3354(69443)

what is the ratio of the circumference of the circle to the area of the circle?
----------
Find the diagonal of the square.
r = d/2
----
C = 2pi*r
A = pi*r^2
----
2pi*r/(pi*r^2) = 2/r